An angle measures 38 degrees less than its supplement. How many degrees are in the angle?
Answer by stanbon(75887)

An angle measures 38 degrees less than its supplement.
How many degrees are in the angle?
---
Let the angle be "x".
Its supplement is "180-x".
--------------------------------
Equation:
supp - angle = 38 degrees
180-x -x = 38
-2x = --142
x = 71 degrees
=================
Cheers,
Stan H.